OpenStudy (shelby1290):

Solve each system of linear equations by elimination. 3a - 2b + 4 = 0 2a - 5b - 1 = 0

OpenStudy (sleepyjess):

You have to multiply one equation or both to make something cancel out

OpenStudy (shelby1290):

What do i multiply by

OpenStudy (sleepyjess):

I would start with multiplying the second equation by -3

OpenStudy (alexandervonhumboldt2):

3a - 2b + 4 = 0 2a - 5b - 1 = 0

OpenStudy (alexandervonhumboldt2):

move -1 and 4 on right side

OpenStudy (alexandervonhumboldt2):

3a - 2b =-4 2a - 5b =1

OpenStudy (alexandervonhumboldt2):

subctarct equation 2 from euqation 1. this will help 3a-2b-2a+5b=-5

OpenStudy (alexandervonhumboldt2):

a+3b=-5

OpenStudy (alexandervonhumboldt2):

a=-5-3b

OpenStudy (alexandervonhumboldt2):

now substitue a into one of the eqautions and solve for b

OpenStudy (alexandervonhumboldt2):

3( -5-3b)- 2b =-4

OpenStudy (alexandervonhumboldt2):

-15-9b-2b=-4

OpenStudy (alexandervonhumboldt2):

-11b=11

OpenStudy (alexandervonhumboldt2):

b=-11/11=???????????? after finding b substute b into one of the original eqautions and solve for a

OpenStudy (alexandervonhumboldt2):

@shelby1290 tell me what are the values of b and a use my instructions

OpenStudy (shelby1290):

@AlexandervonHumboldt2 don't I have to multiply the second equation by -3 first. You know, like what jess said earlier

OpenStudy (shelby1290):

ok

OpenStudy (sleepyjess):

She has to solve it by elimination.

OpenStudy (sleepyjess):

No, you were trying to do substitution.

OpenStudy (alexandervonhumboldt2):

ahhh well

OpenStudy (shelby1290):

Yea, i need to solve it by elimination^

OpenStudy (alexandervonhumboldt2):

you are both wrong

OpenStudy (alexandervonhumboldt2):

eliminationDemonstrates how to solve a linear system using the technique of addition and subtracting i gogloed it

OpenStudy (alexandervonhumboldt2):

hmm

OpenStudy (sleepyjess):

Yes, and you were doing substitution...

OpenStudy (shelby1290):

I'm so confused right now....

OpenStudy (sleepyjess):

@shelby1290 , fan me please

OpenStudy (shelby1290):

I need to eliminate the systems. Just eliminate.

OpenStudy (alexandervonhumboldt2):

brb fan me too then

OpenStudy (alexandervonhumboldt2):

so ok

Nnesha (nnesha):

 lets do elimination method